OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the second and third equations, then the first and second equations. 5x + 2y + z = -2 3x + 4y + 3z = 2 -4x - 3y - 3z = 1

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

cancel out z in 2nd and 3rd equation

OpenStudy (phi):

after the variable z is eliminated in the second and third equations, that means look at only the 2nd and 3rd equations: 3x + 4y + 3z = 2 -4x - 3y - 3z = 1 eliminate the z means make the numbers in front of z be "equal and opposite" i.e. one the negative of the other. once you get that, add the two equations, and the z term will become zero. In this case, we already have 3 and -3, so just add. what do you get ?

OpenStudy (anonymous):

0

OpenStudy (phi):

add the two equations. do you know how to do that ?

OpenStudy (anonymous):

kind of i will try

OpenStudy (anonymous):

-x+7y=3

OpenStudy (anonymous):

phi you here

OpenStudy (phi):

4y-3y is not 7y

OpenStudy (anonymous):

y

OpenStudy (anonymous):

-x+y=3

OpenStudy (phi):

Is that enough to eliminate some of the choices?

OpenStudy (anonymous):

yes thank you, i have just one more

OpenStudy (phi):

then the first and second equations. 5x + 2y + z = -2 3x + 4y + 3z = 2 In case you had to do this: multiply the first eq by -3 (that will make the # in front of z be -3 which is the opposite of +3 in the 2nd equation: -15x -6y -3z= +6 3x + 4y + 3z = 2 add the 2 equations, just like before.